4) Define the function f :Rx R + RxR by f(x, y) = (x + y, x – 2y). Prove that f is a bijection.
Answer #1

f(1,9) = f(u, v) => (at y, a-2y) = (u+v, U-2v) .. 2+= utvå 2-24 = 6-2V : 22+27 + 2-2y = 2u+2v+U-2V - a= u & y=V So; f(x) = f(